# Deep Learning Framework for History Matching CO2 Storage with 4D Seismic and Monitoring Well Data

### v1

@@ 1 Introduction
Geological carbon storage has the potential to reduce substantially the amount of CO2 emitted to the atmosphere. The ability to predict the performance of injection operations (e.g., pressure buildup, plume location) is essential for the gigatonne-scale deployment of this technology. Flow prediction is challenging, however, in large part because of the high degree of geological uncertainty associated with storage formations. Data assimilation procedures, often referred to as history matching in this setting, are typically applied in subsurface flow modeling to calibrate the geomodel based on observations. Our goal in this work is to introduce a new deep-learning-based methodology able to efficiently perform data assimilation in situations with a high level of prior uncertainty, and measurements in the form of in-situ (downhole) monitoring data and 4D (time-lapse) seismic data. Data of these types are expected to be available in large-scale carbon storage operations. Our focus here is on early-time predictions, when uncertainty is high but performance assessment is essential.
@@
Both carbon storage modeling and data assimilation for subsurface operations have been considered extensively in the literature. Reviews of carbon storage, including discussion of challenges, current operations, and modeling strategies, have been provided by Aminu et al. (2017) and Eigbe et al. (2023). Our review here will focus on data assimilation and its use in geological carbon storage. For general background on history matching, please see Oliver et al. (2008). History matching algorithms can be basically divided into gradient-based and derivative-free methods (Rwechungura et al., 2011). Ensemble-based methods, which are a type of derivative-free method, are now widely applied. These include the well known ensemble Kalman filter (EnKF) (Evensen et al., 2007) and ensemble smoother with multiple data assimilation (ESMDA) (Emerick and Reynolds, 2013) algorithms, among many others.
@@
The introduction of different types of surrogate models, applied in place of computationally intensive forward simulation models, has enabled the use of more rigorous (yet more demanding) history matching methods. For example, Chen et al. (2018) constructed a multivariate adaptive regression spline (MARS) proxy for use in CO2 storage modeling and data assimilation. Rana et al. (2018) applied a Gaussian-process-based surrogate model to history match a coalbed methane reservoir. Wang et al. (2021) utilized a differentiable neural network surrogate model to obtain gradient information for use in history matching.
@@
The development of deep learning models for subsurface flow has been a very active research area, and a wide range of network architectures have been considered. These include convolutional neural networks (CNN) (Mo et al., 2019), recurrent neural networks (RNN) (Zhang et al., 2023), graph neural networks (GNN) (Tang and Durlofsky, 2024), and Fourier neural operators (FNO) (Wen et al., 2023), among many others. Deep learning surrogates have been extensively explored for history matching of subsurface flow problems. Tang et al. (2020), for example, constructed a ‘recurrent R-U-Net’ surrogate model based on a residual U-Net and a long short term memory (LSTM) network for 2D oil-water flow problems, which they used to history match channelized systems. Han et al. (2024) extended this model to treat CO2 storage problems with multiple geological scenarios. They combined the surrogate model with a hierarchical Markov chain Monte Carlo (MCMC) method to estimate posterior scenario parameters (metaparameters) and realizations. Wang et al. (2022) incorporated domain knowledge into CNNs to construct theory-guided CNN (TgCNN) surrogates for two-phase subsurface flow problems, which they combined with an iterative ensemble smoother for history matching. Zhang et al. (2023) constructed an LSTM-based surrogate model to predict the oil recovery factor. This was combined with Bayesian MCMC for history matching.
@@
In this study, we introduce a new deep-learning-based framework for early-time history matching of geological carbon storage with 4D seismic and monitoring well data. Rather than construct and train a single complex network able to predict both types of data, we develop two separate fit-for-purpose deep learning models. These models are each relatively straightforward to develop and train, which represents a major advantage of our approach. Specifically, for seismic predictions we introduce a 3D U-Net, which takes as input the highly resolved geomodel and provides as output saturation fields at the scale informed by the seismic data. Monitoring well data are, by contrast, highly resolved vertically but very local areally. These data are predicted using a 1D U-Net with multiple channels corresponding to time steps. The training data for both surrogates are obtained from 3500 flow simulations, with the coarser-scale interpreted seismic data generated by filtering the high-resolution saturation fields. The two trained surrogates are incorporated within a hierarchical MCMC method (Han et al., 2024) for history matching. The geomodels are characterized by a set of metaparameters (which include mean and variance of log-permeability), with each set corresponding to an infinite number of realizations. The impact of the seismic data on posterior geomodel metaparameters and the predicted CO2 plume is assessed. This level of analysis would not be feasible without efficient and robust surrogate models.
@@ 2 Model description and interpreted seismic data
In this section we describe the simulation setup and the synthetic geomodels considered in this study. We then discuss how the interpreted 4D seismic data used in our history matching framework are generated from high-fidelity models.
@@ 2.1 Basic modeling setup
In practical CO2 storage operations, the early-time dynamics (e.g., from a few months to 1-2 years) will be of great interest. This is because the aquifer response over this time frame can be used to assess whether the operation is proceeding as expected, to better understand formation properties, and to determine future injection and monitoring strategies. The geomodels considered in this study are designed to capture these early-time dynamics. The models are gridded such that the near-injector flow field is well resolved. The storage aquifer domain is, however, relatively small since it is not our intention to use these models for long-time predictions.
@@
A fully-penetrating vertical injection well is placed near the center of the storage aquifer, at $i=65$, $j=65$ (where $(i,j)$ denotes areal location on the grid), as shown in Figure 1(b). This well is specified to inject 0.5 Mt/year, for a period of one year (as noted earlier, our interest here is in early time behavior). A monitoring well, shown in Figure 1(b), is placed at $i=75$, $j=75$, which is about 100 m away from the injector. The monitoring well here is assumed to provide CO2 saturation at each layer at a set of time steps. Pressure data, if available, could also be included in our framework.
@@ 2.3 Generation of interpreted 4D seismic data
In our framework, rather than use (and invert) ‘raw’ 4D seismic data, we assume that we are provided with estimates of gas (supercritical CO2) saturation. These interpreted estimates are considered to be at the level of resolution achieved by 4D seismic data. This is, we believe, a reasonable approach, since the time-lapse seismic data are often treated and interpreted in a workflow that is distinct from the flow-based data assimilation process. An analogous approach, in the context of history matching for oil reservoir simulation, was applied by Bukshtynov et al. (2015).
@@
The method described above is straightforward and provides us with a means to generate interpreted seismic saturation data, at the appropriate scale, from a high-fidelity geomodel. It is important to note that our treatment for generating interpreted seismic saturation data can be readily modified to include different smoothing, filtering or averaging procedures, or to account for saturation thresholding, etc. Such modifications might be appropriate in some settings to provide improved mappings from high-fidelity saturation fields to interpreted seismic fields. Note also that noise, which is intended to represent data and model errors, will be incorporated into the interpreted seismic data used for history matching. Our specific treatment for this is described in Section 6.1.
Figure 2: Schematic diagram showing the workflow for generating interpreted seismic data from high-fidelity simulation results. Images on the left are $x$-$z$ cross sections from the full 3D model, on the right are $y$-$z$ cross sections from the 3D model.
@@ 3.1 Surrogate model for seismic data
where $N$ denotes the number of geomodel realizations used for training, $\textbf{S}_{\text{seis},i}$ denotes the reference interpreted seismic saturation field for training sample $i$, and $\hat{{\mathbf{S}}}_{\text{seis},i}$ indicates the surrogate model prediction for sample $i$. The model is trained by adjusting the network parameters $\theta_{seis}$ to minimize $L_{\text{seis}}$. This is accomplished with the Adam optimization algorithm (Kingma and Ba, 2014) on a Nvidia A100 GPU. The initial learning rate is set to be $10^{-4}$, which is updated by multiplying by a factor of 0.2 when the performance does not improve for 10 epochs. The minimum learning rate is set to be $10^{-7}$.
@@ 3.2 Surrogate model for monitoring-well data
The two trained surrogate models can now be used for history matching CO2 storage operations where both 4D seismic and monitoring well data are available. Our framework is flexible, and could be easily extended to handle, e.g., pressure data measured at the monitoring well (though this was not attempted here). Either surrogate model could also be modified or enhanced as necessary to treat more complicated geomodels. We now describe the MCMC history matching procedure.
@@ 4 Hierarchical MCMC method and overall workflow
Our data assimilation problem is similar to that considered by Han et al. (2024) in that both metaparameters and associated geomodel realizations must be determined (the previous study did not consider 4D seismic data so the setup is otherwise quite different). Thus, we apply the hierarchical Markov Chain Monte Carlo (MCMC) method used in that work. MCMC is a widely used history matching technique in which geomodels are sampled iteratively and the likelihood, which involves the data mismatch, is evaluated for each sample. After a large number of iterations, the accepted samples can approximate the posterior distributions. In hierarchical MCMC, two levels of sampling are applied – one for the metaparameters $\mathbf{h}$, and one for the PCA latent variables, $\boldsymbol{\xi}$. Upon convergence (termination), we have an estimate of the posterior probability density function (PDF) $p(\boldsymbol{\xi},\mathbf{h}\mid\mathbf{d}_{\text{obs}})$, where $\mathbf{d}_{\text{obs}}$ denotes the observed data. Because MCMC methods require a large number of sequential function evaluations (e.g., $O(10^{4}-10^{5})$, the use of surrogate models is essential in cases with expensive function evaluations, such as ours.
@@ 5.2 Surrogate performance evaluation
We next present time-lapse results for a specific geomodel (Realization 2 in Figures 7 and 8). Interpreted saturation fields at five time steps are displayed in Figure 9. Although slight differences are evident at some steps, it is clear that the general time evolution is captured accurately by the surrogate model.
@@
5.2.2 Monitoring data surrogate evaluation
The surrogate model for monitoring well data, described in Section 3.2, is trained to predict vertically resolved CO2 saturation at 16 time steps (time steps 0, 2, 4, 6, $\ldots$, 30), one of which is the initial condition. The network contains 16 output channels, with each channel predicting saturation at a particular time step. For training, the batch size is set to 10, the model is trained for 150 epochs, and the initial learning rate is 0.001. This model requires only about 15 minutes for training on a Nvidia A100 GPU. Training is very fast in this case because this surrogate uses only local geomodel properties as inputs rather than the entire field, and it predicts saturations only for a single column of grid blocks (at 16 time steps).

### v2

@@ 1 Introduction
Geological carbon storage has the potential to reduce substantially the amount of CO2 emitted to the atmosphere. The ability to predict the performance of injection operations (e.g., pressure buildup, plume location) is essential for the gigatonne-scale deployment of this technology. Flow prediction is challenging, however, in large part because of the high degree of geological uncertainty associated with storage formations. Data assimilation procedures, often referred to as history matching in this setting, are typically applied in subsurface flow modeling to calibrate the geomodel based on observations. Our goal in this work is to introduce a new deep-learning-based methodology able to efficiently perform data assimilation in situations with a high level of prior uncertainty, and measurements in the form of in-situ (downhole) monitoring data and 4D (time-lapse) seismic data. Data of these types are expected to be available in large-scale carbon storage operations. Our focus here is on early-time predictions, when uncertainty is high but performance assessment is essential, though our methodology is applicable in any phase of a storage project.
@@
Both carbon storage modeling and data assimilation for subsurface operations have been considered extensively in the literature. Reviews of carbon storage, including discussion of challenges, current operations, and modeling strategies, have been provided by Aminu et al. (2017) and Eigbe et al. (2023). Our review here will focus on history matching. History matching algorithms can be divided into gradient-based and derivative-free methods (Rwechungura et al., 2011). Gradient-based methods, like steepest descent method (Chavent et al., 1975), Gauss-Newton method (Li et al., 2003), conjugate gradient method (Lee et al., 1986), etc., often converge quickly, but they require the computation of gradient information, which can be cumbersome or expensive. In addition, these methods essentially search locally, and may therefore not provide a fully representative set of posterior models. Derivative-free methods comprise a broad category of approaches. These include ensemble-based methods, e.g., the well-known ensemble Kalman filter (EnKF) (Evensen et al., 2007; Zovi et al., 2017) and ensemble smoother with multiple data assimilation (ESMDA) (Emerick and Reynolds, 2013; Todaro et al., 2021) algorithms, which are widely applied. These methods are efficient, though they entail some underlying assumptions of Gaussianity that may limit their applicability. Markov chain Monte Carlo (MCMC) methods are another type of derivative-free method that have also been used for history matching (Oliver et al., 1997; Han et al., 2024). These methods are more general as they involve fewer underlying assumptions, though they may require large numbers of function evaluations to achieve convergence.
@@
Different types of surrogate models, applied in place of computationally intensive forward simulation runs, have been used with many different history matching methods. For example, Chen et al. (2018a) constructed a multivariate adaptive regression spline (MARS) proxy for use in CO2 storage modeling and data assimilation. Chen et al. (2020) constructed a bagging MARS (BMARS) surrogate for calibration of a CO2 reservoir model. Rana et al. (2018) applied a Gaussian-process-based surrogate model to history match a coalbed methane reservoir. Wang et al. (2021) utilized a differentiable neural network surrogate model to obtain gradient information for use in history matching. Aslam et al. (2024) introduced a reduced-order model (ROM) to accelerate history matching of fractured simulation models. In this approach, the fractures are represented by a CGNet-Frac model. The ROM was history matched to well production data by adjusting model parameters using a gradient-based optimization method.
@@
The development of deep learning models for subsurface flow has been a very active research area, and a wide range of network architectures have been considered. These include convolutional neural networks (CNN) (Mo et al., 2019), recurrent neural networks (RNN) (Zhang et al., 2023), graph neural networks (GNN) (Tang and Durlofsky, 2024b), and Fourier neural operators (FNO) (Yan et al., 2022; Tang et al., 2024), among many others. Deep learning surrogates have been extensively explored for history matching of subsurface flow problems. Tang et al. (2020), for example, constructed a recurrent R-U-Net surrogate model based on a residual U-Net and a long short term memory (LSTM) network for 2D oil-water flow problems, which they used to history match channelized systems. Han et al. (2024) extended this model to treat CO2 storage problems with multiple geological scenarios. They combined the surrogate model with a hierarchical Markov chain Monte Carlo (MCMC) method to estimate posterior scenario parameters (metaparameters) and realizations. Wang et al. (2022) incorporated domain knowledge into CNNs to construct theory-guided CNN (TgCNN) surrogates for two-phase subsurface flow problems, which they combined with an iterative ensemble smoother for history matching. Zhang et al. (2023) constructed an LSTM-based surrogate model to predict the oil recovery factor. This was combined with Bayesian MCMC for history matching.
@@
In this study, we introduce a new deep learning-based framework for history matching of geological carbon storage with 4D seismic and monitoring well data. Our framework includes both new and existing components, though the integration of these various elements and their effective use for history matching with widely different data types is entirely new. In our procedure, rather than construct and train a single complex network to predict both types of data, we introduce two separate fit-for-purpose deep learning models. These models are each relatively straightforward to develop and train, which represents a major advantage of our approach. In particular, the total training time for both networks is about 2.25 hours. This represents a high level of efficiency when contrasted with training times of tens of hours or more for CNN-RNN or GNN procedures (Han et al., 2024; Tang and Durlofsky, 2024a). Another new treatment introduced here is the scale change applied in the seismic surrogate model. Specifically, the 3D U-Net developed in this work takes as input the highly resolved geomodel, though the saturation output it provides is at a different scale; namely, the scale informed by the seismic data. Monitoring well data, which are highly resolved vertically but very local areally, are predicted using a 1D U-Net with multiple channels corresponding to time steps.
The two trained surrogates are incorporated within a hierarchical MCMC method (Han et al., 2024) for history matching. Higher degrees of prior uncertainty are considered here than in many previous history matching studies. In particular, our geomodels are uncertain both in their metaparameters (e.g., mean and variance of log-permeability, permeability anisotropy ratio) and in terms of detailed heterogeneous realizations. For these highly uncertain systems, we assess the impact of the seismic data on uncertainty reduction and on the predicted CO2 plume using two different treatments. In one case, the seismic data are assumed to provide an estimate of the CO2 saturation value (which may be possible in some practical settings), while in the other case the seismic data indicate only the presence or absence of CO2. The data assimilation capabilities and detailed results presented here would not be feasible without the efficient and robust workflow developed in this study.
@@ 2 Model description and interpreted seismic data
In this section, we describe the simulation setup and the synthetic geomodels considered in this study. We then discuss how the interpreted 4D seismic data used in our history matching framework are generated from high-fidelity models.
@@ 2.1 Basic modeling setup
In practical CO2 storage operations, the early-time dynamics (e.g., from a few months to 1-2 years) will be of great interest. This is because the aquifer response over this time frame can be used to assess whether the operation is proceeding as expected, to better understand formation properties, and to determine future injection and monitoring strategies. The geomodels considered in this study are designed to capture these early-time dynamics. The models are gridded such that the near-injector flow field is well resolved. The storage aquifer domain is relatively small since it is not our intention to use the model for long-time predictions.
@@
Although our immediate interest is in predictions over a one-year time frame, the methodology introduced in this study is also directly applicable for longer-term predictions (years to decades). In such cases, the CO2 plume will extend farther, which means that the high-resolution central storage aquifer domain will need to be larger (areally) than 896 m $\times$ 896 m. This in turn will require the use of more grid cells in this region, leading to higher computational costs for the full-order training simulations and network training.
A fully penetrating vertical injection well is placed near the center of the storage aquifer, at $i=65$, $j=65$ (where $(i,j)$ denotes areal location on the grid), as shown in Figure 1(b). This well is specified to inject 0.5 Mt/year, for a period of one year. A monitoring well, shown in Figure 1(b), is placed at $i=75$, $j=75$, which is about 100 m away from the injector. The monitoring well here is assumed to provide CO2 saturation in each layer at a set of time steps. Pressure data, if available, could also be included in our framework.
@@ 2.2 Metaparameters and geomodels
In this work, the dimension of $\boldsymbol{\xi}$ ($n_{d}$) is set to 800, which preserves 89% of the ‘energy’ (sum of squared singular values) of the original realizations. The use of $n_{d}=800$ is sufficient to retain the key features and the important fine-scale detail in the SGeMS models. This is illustrated in Figure 2, where we display an SGeMS realization and the corresponding PCA reconstruction. The use of a lower value of $n_{d}$ could result in an overly smoothed PCA model, but this is not observed with $n_{d}=800$.
(a) Original SGeMS field
(b) PCA reconstruction with $n_{d}=800$
Figure 2: Original SGeMS and reconstructed PCA fields.
@@ 2.3 Generation of interpreted 4D seismic data
In our framework, rather than use (and invert) ‘raw’ 4D seismic data, we assume that we are provided with estimates of gas (supercritical CO2) saturation or plume location (by plume location we mean the presence of gas above some threshold saturation). These interpreted estimates are considered to be at the level of resolution achieved by 4D seismic data. This is, we believe, a reasonable approach, since the time-lapse seismic data are often treated and interpreted in a workflow that is distinct from the flow-based data assimilation process. An analogous approach, in the context of history matching for oil reservoir simulation, was applied by Bukshtynov et al. (2015).
@@
In practice, seismic data have been widely used for monitoring CO2 plume migration  (Chadwick et al., 2005; Zhang et al., 2012). Ideally, the seismic data can be inverted to provide an estimate of CO2 saturation ($S_{g}$), using full-waveform inversion (FWI) (Dupuy et al., 2017; Queißer and Singh, 2013) or even deep learning models (Um et al., 2023). However, in some cases, the seismic data may be informative only in terms of the presence or absence of the plume. We refer to this as plume location identification. Plume location is represented in our modeling with a value of 1 where the plume is present (meaning $S_{g}$ exceeds a detection limit or threshold value), and a value of 0 otherwise. A binary image, as shown in Figure 3, then indicates the presence or absence of the plume. In this work the threshold value is set to 0.05, which is a reasonable practical limit. The use of both types of interpreted seismic data, i.e., seismic-scale $S_{g}$ fields and binary plume location, will be considered separately in our history matching.
The treatments described above are straightforward and provide us with a means to generate interpreted seismic saturation data or plume location data, at the appropriate scale, from a high-fidelity geomodel. It is important to note that our approaches for generating interpreted seismic data can be readily modified to include different smoothing, filtering or averaging procedures, etc. It is also straightforward to consider different threshold values in the plume location approach. In either case, the treatment should be consistent with the actual seismic data acquisition and processing workflow. Note also that noise, intended to represent data and model errors, is incorporated into the interpreted seismic data during history matching. Our specific treatment for this will be described in Section 6.1.
Figure 3: Schematic showing the workflow for generating interpreted seismic data from high-fidelity simulation results. Images on the left are $x$-$z$ cross sections from the full 3D model and those on the right are $y$-$z$ cross sections. Interpreted seismic data can provide estimates of either seismic-scale $S_{g}$ or plume location.
@@ 3.1 Surrogate model for seismic data
where $N$ denotes the number of geomodel realizations used for training, $\textbf{S}_{\text{seis},i}$ denotes the reference interpreted seismic saturation field for training sample $i$, and $\hat{{\mathbf{S}}}_{\text{seis},i}$ indicates the surrogate model prediction for sample $i$. The model is trained by adjusting the network parameters $\theta_{seis}$ to minimize $L_{\text{seis}}$. This is accomplished with the Adam optimization algorithm (Kingma and Ba, 2014) on a Nvidia A100 GPU. The initial learning rate is set to be $10^{-4}$, which is updated by multiplying by a factor of 0.2 when the performance does not improve for 10 epochs. The minimum learning rate is set to be $10^{-7}$. When plume location rather than saturation is needed, the predicted seismic saturation field is thresholded to provide the binary plume location field (i.e., the plume is considered to be present when $S_{g}>0.05$). Therefore, the same surrogate is used for both seismic saturation and plume location.
@@ 3.2 Surrogate model for monitoring-well data
The two trained surrogate models can now be used for history matching CO2 storage operations where both 4D seismic and monitoring well data are available. Our framework is flexible, and could be easily extended to handle, e.g., pressure data measured at the monitoring well (though this was not considered here). Either surrogate model could also be modified or enhanced as necessary to treat more complicated geomodels. We now describe the MCMC history matching procedure.
@@ 4 Hierarchical MCMC method and overall workflow
Our data assimilation problem is similar to that considered by Han et al. (2024) in that both metaparameters and associated geomodel realizations must be determined (the previous study did not consider 4D seismic data so the setup is otherwise quite different). Thus, we apply the hierarchical Markov Chain Monte Carlo (MCMC) method used in that work. MCMC is a widely used history matching technique in which geomodels are sampled iteratively and the likelihood, which involves the data mismatch, is evaluated for each sample. After a large number of iterations, the accepted samples can approximate the posterior distributions. In hierarchical MCMC, two levels of sampling are applied – one for the metaparameters $\mathbf{h}$, and one for the PCA latent variables $\boldsymbol{\xi}$. Upon convergence (termination), we have an estimate of the posterior probability density function (PDF) $p(\boldsymbol{\xi},\mathbf{h}\mid\mathbf{d}_{\text{obs}})$, where $\mathbf{d}_{\text{obs}}$ denotes the observed data. Because MCMC methods require a large number of sequential function evaluations (e.g., $O(10^{4}-10^{5})$, the use of surrogate models is essential in cases with expensive function evaluations, such as ours.
@@
Before presenting numerical results, it is useful to discuss some important differences between hierarchical MCMC and other history matching methods. In practical settings, ensemble-based methods are commonly applied. ESMDA (Emerick and Reynolds, 2013), for example, was used by Crain et al. (2024) to history match a CO2 storage project in the Illinois Storage Corridor in the US. Ensemble-based methods, however, typically involve several limiting assumptions that are not compatible with our setup. Specifically, these methods are generally used for cases with fixed metaparameters. Even if they are applied hierarchically to consider metaparameter uncertainty, these methods still require Gaussian priors (Oliver, 2022). The setup here is more general in that we consider a range of metaparameters that can have any prior distribution (uniform rather than Gaussian priors are used in our examples). Thus, standard implementations of ensemble-based methods are not applicable, so direct numerical comparisons with such approaches are not straightforward.
Standard (rather than hierarchical) MCMC methods are also difficult to compare against because the dimension of the geomodels is very high, even using a dimension reduction procedure such as PCA. Thus, it may be difficult for the Markov chain to converge in a standard approach. Our hierarchical MCMC procedure is ‘dimension robust’ (Chen et al., 2018b), meaning the sampling efficiency does not degrade as the dimension of the PCA representation increases. This allows our method to achieve high-dimensional sampling more efficiently than in a standard MCMC approach.
@@ 5.2 Surrogate performance evaluation
We next present time-lapse results for a specific geomodel (Realization 2 in Figures 8 and 9). Interpreted saturation fields at five time steps are displayed in Figure 10. Although slight differences are evident at some steps, it is clear that the general time evolution is captured accurately by the surrogate model.
@@
5.2.2 Monitoring well data surrogate evaluation
The surrogate model for monitoring well data, described in Section 3.2, is trained to predict vertically resolved CO2 saturation at 16 time steps (time steps 0, 2, 4, 6, $\ldots$, 30). The first time step corresponds to the initial condition. The network contains 16 output channels, with each channel predicting saturation at a particular time step. For training, the batch size is set to 10, the model is trained for 150 epochs, and the initial learning rate is 0.001. This model requires only about 15 minutes for training on a Nvidia A100 GPU. Training is very fast in this case because this surrogate uses only local geomodel properties as inputs rather than the entire field, and it predicts saturations only for a single column of grid blocks (at 16 time steps).